# CoTKR: Chain-of-Thought Enhanced Knowledge Rewriting for Complex Knowledge Graph Question Answering

### v3

@@ A.2 Datasets
GrailQA [^5] This dataset is distributed under the CC BY-SA 4.0 license and our utilization complies with the terms specified in the license. Gu et al. (2021) is a challenging, large-scale multi-hop KGQA benchmark. It is an English dataset that utilizes Freebase Bollacker et al. (2007, 2008) as KG. It spans 86 domains, such as Sports, Location, and Computer Video Games, and comprises 64,331 questions (44,337 train, 6,763 dev, 13,231 test). This dataset features a large number of entities and relations, complex logical forms, and noise in entity mentions within the questions. The training and dev sets provide annotated SPARQL queries and answer entities, while the test set comprises only the questions. For evaluation convenience, the dev set is used for testing.
GraphQuestions [^6] This dataset is licensed under the Creative Commons Attribution 4.0 and our usage aligns with the intended purposes outlined in this license. Su et al. (2016) is a characteristic-rich dataset for factoid question answering based on Freebase across 70 domains, like People, Astronomy, and Medicine. This English dataset focuses on the following question characteristics: structure complexity, function, commonness, paraphrasing, and answer cardinality. It comprises 5,166 questions (2,771 train, 2,395 test), with nearly half requiring multi-hop reasoning. For each question, the dataset provides corresponding SPARQL query and answer entities.
@@ A.3 Large Language Models
Llama-2777The license of Llama-2 is available at https://ai.meta.com/resources/models-and-libraries/llama-downloads/. Touvron et al. (2023b), an updated version of Llama-1 Touvron et al. (2023a), is developed using a training corpus comprising 2 trillion tokens and features a context length twice that of Llama-1. To better accomplish the knowledge rewriting task, we select Llama-2-7B-Chat [^8] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
Llama-3999The license of Llama-3 is available at https://llama.meta.com/llama3/license. AI@Meta (2024) is the latest model in the Llama series. It is renowned for its mastery of language nuances, contextual comprehension, and proficiency in executing complex tasks such as translation and generating dialogues. We choose Llama-3-8B-Instruct [^10] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for knowledge rewriting.
Mistral [^11] It is under the Apache 2 License. Jiang et al. (2023a) is an open-source LLM developed by Mistral AI. We select the latest instruction-tuned version, Mistral-7B-Instruct-v0.3121212https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, as our QA model.
@@ A.4 Retrieval Methods Details
2-Hop subgraph is a naïve question-related context. Most KBQA studies under RAG paradigms consider triples within the N-hop subgraph of the head entity as contextual knowledge Baek et al. (2023); Sen et al. (2023); Wang et al. (2024); Ko et al. (2024). To retrieve the 2-hop subgraph around the head entity, we execute SPARQL queries on Freebase. Given the large size of the 2-hop subgraph, we use all-MiniLM-L6-v2151515https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode all 1-hop and 2-hop relations of the head entity and the question, excluding meaningless relations, such as “common.topic.webpage”. Then, we select semantically similar relation paths based on cosine similarity. Finally, we sample the corresponding triples from KG based on these relation paths. In this experiment, we select the top 30 triples as our retrieval results. However, the small size of the 2-hop subgraphs for some entities may result in fewer than 30 triples being retrieved.
BM25 Robertson and Zaragoza (2009) is a retrieval method based on TF-IDF scores of sparse word matching between input questions and passages. This method is commonly used for integrating multimodal data sources or text-based QA Wang et al. (2023a); Yu et al. (2023). We follow the processing method in DecAF Yu et al. (2023), simply linearizing the 1-hop subgraph of the topic entity as the article. We use the BM25 implemented by Pyserini Lin et al. (2021) and collect the triples corresponding to the candidate articles as the retrieval result. Specifically, we initially retrieve 10 candidate articles, each containing up to 10 triples. Subsequently, we remove any redundant triples or those containing meaningless relations. Given the limited context length of LLMs, we select the top 30 triples as the context information for question answering. After filtering, the number of triples in the candidate articles may be less than 30, thus resulting in the retrieval subgraphs for some questions containing fewer than 30 triples.
@@ A.5 Implementation Details
We utilize LoRA Hu et al. (2022) to achieve parameter-efficient fine-tuning. For supervised fine-tuning and DPO, the batch size, learning rate, lora rank, lora alpha, and lora dropout are set to 128, 1e-4, 64, 128, and 0.05, respectively. In supervised fine-tuning, we train for 10 epochs and save the best model based on validation set results. In DPO, we observe that more training steps may lead to decreased model performance. Consequently, we train for 1 to 2 epochs on GraphQuestions using approximately 2,000 training samples, and for 5 to 20 steps on GrailQA using 640 to 2560 training samples. During inference, the temperature is set to 0 for ChatGPT and 0.01 for open-source LLMs, ensuring relatively stable output. All the parameter settings mentioned above are the optimal results after multiple trials.
@@
[…]
@@
[…]
@@ C.1 Experiments on ComplexWebQuestions
To evaluate the robustness of CoTKR, we conduct our experiments on ComplexWebQuestionsTalmor and Berant (2018). We utilize ChatGPT as the knowledge rewriter, Mistral as the question-answering model, and 2-Hop as the retrieval method. The experimental results, presented in Table 9, demonstrate the effectiveness of CoTKR.
@@ C.4 GPT-4-score as Evaluation Metrics
Given the powerful natural language understanding and generation capabilities of closed-source large models, many existing works employ ChatGPT as an evaluator to provide high-quality evaluation results Sottana et al. (2023); Liu et al. (2023); Min et al. (2023). In our approach, we use GPT-4 as the evaluator to assess whether all answer entities are present in the responses. We refer to this evaluation metric as GPT-4-score. Compared to EM, this metric is more flexible, as LLMs are capable of recognizing synonyms of answer entities. We use it to evaluate the first 300 questions from GrailQA using Llama-3 as knowledge rewriter, ChatGPT as question-answering model, and 2 hop as retrieval method. We provide the prompt for GPT-4-score in Table 12.

### v4

@@ A.2 Datasets
GrailQA [^5] This dataset is distributed under the CC BY-SA 4.0 license and our utilization complies with the terms specified in the license. DBLP:conf/www/GuKVSLY021 is a challenging, large-scale multi-hop KGQA benchmark. It is an English dataset that utilizes Freebase DBLP:conf/aaai/BollackerCT07; DBLP:conf/sigmod/BollackerEPST08 as KG. It spans 86 domains, such as Sports, Location, and Computer Video Games, and comprises 64,331 questions (44,337 train, 6,763 dev, 13,231 test). This dataset features a large number of entities and relations, complex logical forms, and noise in entity mentions within the questions. The training and dev sets provide annotated SPARQL queries and answer entities, while the test set comprises only the questions. For evaluation convenience, the dev set is used for testing.
GraphQuestions [^6] This dataset is licensed under the Creative Commons Attribution 4.0 and our usage aligns with the intended purposes outlined in this license. DBLP:conf/emnlp/SuSSSGYY16 is a characteristic-rich dataset for factoid question answering based on Freebase across 70 domains, like People, Astronomy, and Medicine. This English dataset focuses on the following question characteristics: structure complexity, function, commonness, paraphrasing, and answer cardinality. It comprises 5,166 questions (2,771 train, 2,395 test), with nearly half requiring multi-hop reasoning. For each question, the dataset provides corresponding SPARQL query and answer entities.
@@ A.3 Large Language Models
Llama-2777The license of Llama-2 is available at https://ai.meta.com/resources/models-and-libraries/llama-downloads/. touvron2023llama2, an updated version of Llama-1 DBLP:journals/corr/abs-2302-13971, is developed using a training corpus comprising 2 trillion tokens and features a context length twice that of Llama-1. To better accomplish the knowledge rewriting task, we select Llama-2-7B-Chat [^8] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
Llama-3999The license of Llama-3 is available at https://llama.meta.com/llama3/license. llama3modelcard is the latest model in the Llama series. It is renowned for its mastery of language nuances, contextual comprehension, and proficiency in executing complex tasks such as translation and generating dialogues. We choose Llama-3-8B-Instruct [^10] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for knowledge rewriting.
Mistral [^11] It is under the Apache 2 License. jiang2023mistral is an open-source LLM developed by Mistral AI. We select the latest instruction-tuned version, Mistral-7B-Instruct-v0.3121212https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, as our QA model.
@@ A.4 Retrieval Methods Details
2-Hop subgraph is a naïve question-related context. Most KBQA studies under RAG paradigms consider triples within the N-hop subgraph of the head entity as contextual knowledge baek-etal-2023-knowledge; sen-etal-2023-knowledge; DBLP:journals/corr/abs-2401-00426; DBLP:journals/corr/abs-2403-02966. To retrieve the 2-hop subgraph around the head entity, we execute SPARQL queries on Freebase. Given the large size of the 2-hop subgraph, we use all-MiniLM-L6-v2151515https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode all 1-hop and 2-hop relations of the head entity and the question, excluding meaningless relations, such as “common.topic.webpage”. Then, we select semantically similar relation paths based on cosine similarity. Finally, we sample the corresponding triples from KG based on these relation paths. In this experiment, we select the top 30 triples as our retrieval results. However, the small size of the 2-hop subgraphs for some entities may result in fewer than 30 triples being retrieved.
BM25 DBLP:journals/ftir/RobertsonZ09 is a retrieval method based on TF-IDF scores of sparse word matching between input questions and passages. This method is commonly used for integrating multimodal data sources or text-based QA DBLP:journals/corr/abs-2308-13259; DBLP:conf/iclr/YuZNZL0HWWX23. We follow the processing method in DecAF DBLP:conf/iclr/YuZNZL0HWWX23, simply linearizing the 1-hop subgraph of the topic entity as the article. We use the BM25 implemented by Pyserini DBLP:conf/sigir/LinMLYPN21 and collect the triples corresponding to the candidate articles as the retrieval result. Specifically, we initially retrieve 10 candidate articles, each containing up to 10 triples. Subsequently, we remove any redundant triples or those containing meaningless relations. Given the limited context length of LLMs, we select the top 30 triples as the context information for question answering. After filtering, the number of triples in the candidate articles may be less than 30, thus resulting in the retrieval subgraphs for some questions containing fewer than 30 triples.
@@ A.5 Implementation Details
We utilize LoRA DBLP:conf/iclr/HuSWALWWC22 to achieve parameter-efficient fine-tuning. For supervised fine-tuning and DPO, the batch size, learning rate, lora rank, lora alpha, and lora dropout are set to 128, 1e-4, 64, 128, and 0.05, respectively. In supervised fine-tuning, we train for 10 epochs and save the best model based on validation set results. In DPO, we observe that more training steps may lead to decreased model performance. Consequently, we train for 1 to 2 epochs on GraphQuestions using approximately 2,000 training samples, and for 5 to 20 steps on GrailQA using 640 to 2560 training samples. During inference, the temperature is set to 0 for ChatGPT and 0.01 for open-source LLMs, ensuring relatively stable output. All the parameter settings mentioned above are the optimal results after multiple trials.
@@
KG-to-Text The International System of Units includes various units for different physical quantities such as length, area, potential, substance, conductance, volume, resistance, current, force, and pressure. For example, the unit for electric field strength in volts per meter is 1.0. Additionally, it covers a wide range of units including square kilometers, square meters The unit of area in the International System of Units that has an electric field strength unit of volt per meter is square meters (m²).
Summary The International System of Units includes various measurement systems for different physical quantities such as length, area, volume, conductance, resistance, current, force, and pressure. The measurement system with an electric field strength unit of Volt per meter does not directly relate to a specific area unit within the International System of Units. There is no direct relationship between the measurement system with an electric field strength unit of Volt per meter and any specific area unit within the International System of Units.
CoTKR Reason: I need to know the unit of area associated with the measurement system that has an electric field strength unit of Volt per meter.
@@
CoTKR+PA Reason: I need to know which measurement system includes the unit for electric field strength in Volt per meter.
@@ C.1 Experiments on ComplexWebQuestions
To evaluate the robustness of CoTKR, we conduct our experiments on ComplexWebQuestionsDBLP:conf/naacl/TalmorB18. We utilize ChatGPT as the knowledge rewriter, Mistral as the question-answering model, and 2-Hop as the retrieval method. The experimental results, presented in Table 9, demonstrate the effectiveness of CoTKR.
@@ C.4 GPT-4-score as Evaluation Metrics
Given the powerful natural language understanding and generation capabilities of closed-source large models, many existing works employ ChatGPT as an evaluator to provide high-quality evaluation results DBLP:conf/emnlp/SottanaLZY23; DBLP:conf/emnlp/LiuIXWXZ23; DBLP:conf/emnlp/MinKLLYKIZH23. In our approach, we use GPT-4 as the evaluator to assess whether all answer entities are present in the responses. We refer to this evaluation metric as GPT-4-score. Compared to EM, this metric is more flexible, as LLMs are capable of recognizing synonyms of answer entities. We use it to evaluate the first 300 questions from GrailQA using Llama-3 as knowledge rewriter, ChatGPT as question-answering model, and 2 hop as retrieval method. We provide the prompt for GPT-4-score in Table 12.